OpenStudy (anonymous):

a cylindrical container of V=450 is to be made of different material for bottom is R60per square metre, for curved side is R200per square metre and the lid is R160per square metre. what is the least cost of the container in rands?

OpenStudy (anonymous):

v = pi * (radius^3) r^3 = 450 / pi radius = \[\sqrt[3]{450/\pi}\]

OpenStudy (anonymous):

v=450 cubic metre

OpenStudy (amistre64):

surface area = pi r^2(R60) + pi r^2(R160) + 2pi r(R200)

OpenStudy (amistre64):

surface area = pi r^2(R60) + pi r^2(R160) + 2pi r(R200)(h) volume = pi r^2 (height) 450 = pi r^2 * h 450/pi r^2 = h ------------------------------ surface area = pi r^2(60) + pi r^2(160) + 2pi r(200)(450/pi r^2)

OpenStudy (amistre64):

Sa = pi r^2(60+160) + 2.200.450/r Sa = pi r^2(220) + 180000/r Sa = (pi r^3(220) + 180000)/r Sa' = [660 pi r^3 - 220pi r^3 - 180000]/r^2 440 pir^3 - 180000 = 0 when..... r =abt. 2.363

OpenStudy (amistre64):

err... r =abt. 5.07

OpenStudy (anonymous):

yip and cost is R 53 268.91

OpenStudy (amistre64):

the derivative is right; just gotta see if i can find a common answer for r lol

OpenStudy (amistre64):

440 pir^3 = 180000 r^3 = 4500/11 pi r = cbrt(4500/11 pi) perhaps?

OpenStudy (amistre64):

r = abt 5.07...thats good lol

OpenStudy (amistre64):

450/[pi (5.07)^2] = h = abt 5.57

OpenStudy (anonymous):

thanks everyone